# Contrastive Disentangling: Fine-grained representation learning through multi-level contrastive learning without class priors

### v1

@@ Abstract
Recent advancements in unsupervised representation learning often leverage class information to enhance feature extraction and clustering performance. However, this reliance on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous. In this paper, we propose Contrastive Disentangling (CD), a simple and effective framework that learns representations without any reliance on class priors. Our framework employs a multi-level contrastive learning strategy that combines instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations. Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations for different samples, (2) the feature-level contrastive loss promotes independence among the feature head predictions, and (3) the normalized entropy loss encourages the feature heads to capture meaningful and prevalent attributes from the data. These components work together to enable CD to significantly outperform existing methods, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10, particularly in scenarios where class priors are absent. The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
@@
Disentangling representations , Clustering , Contrastive learning , Representation learning , Unsupervised learning
@@ 1 Introduction
The growing importance of unsupervised learning in processing large-scale unlabeled datasets has increased the demand for effective feature extraction techniques that do not rely on predefined class labels. In computer vision tasks, unsupervised representation learning has become a powerful solution for deriving meaningful feature representations from raw data. A key advancement in this domain is contrastive learning, which has emerged as a prominent paradigm within unsupervised representation learning. By maximizing consistency between different augmented views of the same data instance while minimizing similarity between different instances, contrastive learning has significantly advanced feature extraction (Ye et al., 2019; Chen et al., 2020).
As research in this field progresses, numerous contrastive learning methods (Van Gansbeke et al., 2020; Li et al., 2021; Qian, 2023) have started to utilize class number information to further improve performance. These approaches use class numbers as prior knowledge during network construction, achieving state-of-the-art results in unsupervised clustering tasks. This suggests that incorporating class information during training yield more discriminative feature representations. However, in practical applications, class information is often unavailable, or the data may lack clear structure, limiting the applicability of these methods.
To address this limitation, we propose a novel framework, Contrastive Disentangling (CD), which enhances feature extraction and disentanglement in unsupervised learning tasks through a multi-level contrastive learning strategy and an innovative disentanglement loss function, all without relying on class number information. Unlike methods that rely on additional complex modules, such as self-labeling, our approach focuses on validating the effectiveness of the core framework to ensure its broad applicability.
Figure 1: Multi-level Contrastive Learning in the CD Framework. Instance-level (top) operates on image instances, while feature-level (bottom) focuses on features visualized with LIME.
As illustrated in Figure 1, the CD framework incorporates two levels of contrastive learning: instance-level and feature-level. Instance-level contrastive learning aims to distinguish different image instances within the feature space, while feature-level contrastive learning seeks to differentiate the features captured by distinct prediction heads. In the feature-level module, we employ the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to interpret the outputs of specific feature prediction heads, providing an intuitive visualization of the diverse feature attributes captured by the model across the image batch. In summary, the main contributions of this paper include:
@@
    We introduce feature prediction heads, a novel addition to the feature learning process. Using normalized entropy loss, these heads effectively capture common features prevalent in the dataset, thereby enhancing the representational capacity and generalization of the features.
@@
    A feature-level contrastive learning strategy is proposed, ensuring that different feature prediction heads focus on learning distinct and complementary features, thus promoting feature disentanglement.
@@
    We propose a simple and easy-to-use framework that significantly enhances the learning of fine-grained features in unsupervised learning tasks without relying on class information. Extensive experiments across multiple benchmark datasets demonstrate the superior performance of the CD framework, particularly in metrics like NMI, where it outperforms compared methods. Through comprehensive ablation studies, we validate the effectiveness of each component, providing new insights in unsupervised representation learning.
@@ 2.1 Unsupervised Representation Learning
Unsupervised representation learning seeks to extract meaningful features from unlabeled data. The field is primarily dominated by two approaches: generative models and contrastive learning. Generative models (Bengio et al., 2006; Kingma, 2013; Goodfellow et al., 2014) learn to generate samples resembling the true data distribution by modeling the underlying latent space. However, these models often struggle to learn invariant feature representations, which are insensitive to certain input transformations like rotation, translation, and scaling (Ntelemis et al., 2023).
To overcome these challenges, instance-level contrastive learning methods (He et al., 2020; Chen et al., 2020; Grill et al., 2020; Chen and He, 2021) have been proposed. These approaches significantly enhance the model’s capacity to extract invariant features by maximizing the similarity between different views of the same instance. These methods have exhibited superior performance across various downstream tasks, underscoring their efficacy in unsupervised learning scenarios.
2.2 Utilization of Class Number Information
Recent approaches have leveraged class number information as additional prior knowledge to enhance model performance. Methods (Van Gansbeke et al., 2020; Li et al., 2021; Niu et al., 2022; Qian, 2023) incorporate class number to constrain the final output during network construction. Several of these methods (Van Gansbeke et al., 2020; Niu et al., 2022; Qian, 2023) also employ pseudo-labeling techniques, further enhancing the model’s performance in unsupervised clustering tasks. These advancements indicate that class-level information can effectively assist models in distilling more semantically meaningful features.
However, this dependence on class information constrains the applicability of these methods in real-world settings where class labels or the number of classes are frequently unknown or ambiguous. This limitation emphasizes the necessity for more adaptive approaches that can perform effectively without relying on class priors.
@@ 2.3 Multi-level Contrastive Learning
The introduction of cluster-level contrastive learning by CC (Li et al., 2021) for clustering tasks has unveiled new avenues in contrastive learning research. However, partitioning based on predefined categories may not adequately capture the fine-grained features inherent in datasets, potentially constraining the model’s feature extraction capabilities. Building upon this observation, we propose a novel architecture that leverages feature space representations as inputs to feature prediction heads. By extracting common features and implementing feature-level contrastive learning, our approach achieves relative decoupling of distinct feature heads. This facilitates the capture of more discriminative and fine-grained features, thereby advancing the state-of-the-art in unsupervised representation learning.
@@ 3.1 Proposed Contrastive Disentangling Framework
The proposed CD framework leverages multi-level contrastive learning, as illustrated in Figure 2. It consists of the following key components:
Figure 2: Architecture of the proposed Contrastive Disentangling (CD) framework. The top section illustrates the shared backbone network, instance-level projector, and feature-level predictor, while the bottom section displays the optimization objectives for instance-level and feature-level contrastive learning.
@@
    Data Augmentation. Each input batch $x$ undergoes a series of data augmentation operations, such as random cropping, horizontal flipping, color distortion, and blurring. These augmentations generate two correlated views, denoted as $x^{(1)}$ and $x^{(2)}$, enabling the model to focus on semantically relevant features that remain invariant under these transformations.
@@
    Feature Extraction via ResNet Backbone. A ResNet architecture (He et al., 2016) is used as the shared backbone network $f(\cdot)$, which extracts feature representations from the augmented views $x^{(1)}$ and $x^{(2)}$. This choice aligns with prevailing practices in contrastive learning (Chen et al., 2020; Li et al., 2021). We denote the extracted feature representations as $h^{(1)}=f(x^{(1)})$ and $h^{(2)}=f(x^{(2)})$.
@@
    Instance Projector. The feature representations $h^{(1)}$ and $h^{(2)}$ are passed through an instance projector $g(\cdot)$, implemented as a multi-layer perceptron (MLP). This MLP consists of two linear layers, batch normalization (Ioffe, 2015), and a ReLU activation function. The output is a set of latent representations $z^{(1)}=g(h^{(1)})$ and $z^{(2)}=g(h^{(2)})$, which are crucial for the contrastive learning process (Chen et al., 2020).
@@
    Feature Predictor. The latent representations $z^{(1)}$ and $z^{(2)}$ are fed into a feature predictor $h(\cdot)$, structured similarly to the instance projector but using sigmoid activation in the final layer. The feature predictor outputs predictions $y^{(1)}=h(z^{(1)})$ and $y^{(2)}=h(z^{(2)})$, organized into multiple prediction heads. Each head predicts a distinct feature attribute. We will validate the design of the feature predictor in Section 4.5.3.
@@
    Loss Functions. The total loss comprises three components: the instance-level contrastive loss, the feature-level contrastive loss, and an additional normalized entropy loss. These losses work together to enhance feature diversity, promote independence among learned features, and ensure a wide range of feature predictions.
@@ 3.2 Instance-level Contrastive Learning
The objective of instance-level contrastive learning is to pull the representations of different augmented views of the same image closer in feature space while pushing apart representations of different images. Specifically, given an input batch $x$, two views $x^{(1)}$ and $x^{(2)}$ are generated via data augmentation, which are then processed by the shared backbone network and instance projector to yield latent representations $z^{(1)}=g(f(x^{(1)}))$ and $z^{(2)}=g(f(x^{(2)}))$.
These latent representations, $z^{(1)}$ and $z^{(2)}$, are concatenated into a matrix $\mathbf{z}\in\mathbb{R}^{2N\times d}$, where $N$ is the batch size and $d$ is the feature dimension:
$\mathbf{z}=\begin{bmatrix}z^{(1)}\\ z^{(2)}\end{bmatrix}=\begin{bmatrix}z_{1}\\ z_{2}\\ \vdots\\ z_{2N}\end{bmatrix}$ (1)
To quantify the similarity between any two latent representations, we introduce the cosine similarity measure $\text{sim}_{z}(i,j)$:
$\text{sim}_{z}(i,j)=\frac{z_{i}\cdot z_{j}}{\|z_{i}\|\|z_{j}\|}$ (2)
where $i,j\in\{1,2,\dots,2N\}$ and $N$ is the batch size. Based on this similarity measure, we optimize the model by maximizing the similarity between different views of the same instance while minimizing the similarity between different instances. The loss function for instance $i$ is defined as:
@@
Here, $\text{pos}(i)$ represents the index of the positive sample corresponding to $i$. If $i\leq N$, then $\text{pos}(i)=i+N$; otherwise, $\text{pos}(i)=i-N$. The temperature parameter $\tau_{\text{inst}}$ within the instance-level contrastive loss governs the smoothness of the similarity distribution. Additionally, the indicator function $\mathbbm{1}_{[k\neq i]}$ assigns a value of 1 when $k\neq i$, thereby excluding the self-similarity of $z_{i}$. The instance-level contrastive loss is then calculated across all instance views:
@@ 3.3 Disentanglement Loss
To enhance the independence and diversity of the learned features, we introduce two complementary loss functions: feature-level contrastive loss and normalized entropy loss. These losses work together to promote the learning of disentangled and fine-grained feature attributes.
Feature-level Contrastive Loss. The feature predictor $h(\cdot)$ is designed with multiple independent prediction heads, each focused on predicting a fine-grained feature attribute. After processing the input views $x^{(1)}$ and $x^{(2)}$ through the backbone network and instance projector to generate latent representations $z^{(1)}$ and $z^{(2)}$, these representations are fed into the feature predictor $h(\cdot)$, yielding the corresponding feature predictions $y^{(1)}$ and $y^{(2)}$.
To align with the instance-level contrastive loss, these predictions $y$ are transposed and organized into a matrix $\mathbf{y}\in\mathbb{R}^{2K\times N}$, where $K$ denotes the number of features and $N$ represents the batch size:
$\mathbf{y}=\begin{bmatrix}y^{(1)\top}\\ y^{(2)\top}\end{bmatrix}=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{2K}\end{bmatrix}$ (5)
In this matrix, $y_{i}$ and $y_{\text{pos}(i)}$ correspond to the prediction results of feature head $i$ on the same batch of samples under two different views. We then define the similarity function $\text{sim}_{y}(i,j)$ to measure the cosine similarity between any two feature prediction vectors:
$\text{sim}_{y}(i,j)=\frac{y_{i}\cdot y_{j}}{\|y_{i}\|\|y_{j}\|}$ (6)
To ensure that feature head $i$ learns features different from other heads, the feature-level contrastive loss is defined as follows:
@@
Here, $\tau_{\text{feat}}$ is the temperature parameter for the feature-level contrastive loss, which controls the degree of similarity enforced between different features. The primary objective is to encourage each prediction head to learn disentangled feature attributes. The final feature-level contrastive loss is calculated as follows:
@@
Normalized Entropy Loss. To further promote the learning of common features, we introduce a normalized entropy loss $\mathcal{L}_{\text{entropy}}$. This loss encourages the model to produce balanced and varied predictions by maximizing the entropy of the predicted outputs. By doing so, it prevents the model from overfitting to specific features and promotes the capture of more generalizable attributes. The normalized binary entropy $H(y_{i})$ for each feature prediction vector $y_{i}$ is defined as:
$H(y_{i})=-\frac{1}{N\log(2)}\sum_{j=1}^{N}\left(y_{ij}\log(y_{ij})+(1-y_{ij})\log(1-y_{ij})\right)$ (9)
Subsequently, the normalized entropy loss $\mathcal{L}_{\text{entropy}}$ is computed as the mean entropy across all feature predictions:
$\mathcal{L}_{\text{entropy}}=\frac{1}{2K}\sum_{i=1}^{2K}H(y_{i})$ (10)
Finally, the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$ combines the feature-level contrastive loss and the normalized entropy loss:
$\mathcal{L}_{\text{disentanglement}}=\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$ (11)
where $\alpha$ is a hyperparameter that controls the weight of the normalized entropy loss. It balances the learning of common features within the dataset and the disentanglement of independent feature representations. Minimizing this combined loss enables our CD framework to effectively facilitate the unsupervised extraction and disentanglement of meaningful features from the input data. The efficacy of $\mathcal{L}_{\text{disentanglement}}$ is further validated in the experimental section 4.5.3.

### v2

# Fine-Grained Representation Learning via Multi-Level Contrastive Learning without Class Priors

@@ Abstract
Recent advances in unsupervised representation learning frequently leverage class information to improve the extraction and clustering of features. However, this dependence on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous. In this paper, we propose Contrastive Disentangling (CD), a simple yet effective framework that learns representations without any relying on class priors. CD employs a multi-level contrastive learning strategy, integrating instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations. Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations between different samples; (2) the feature-level contrastive loss promotes independence among feature prediction heads; and (3) the normalized entropy loss ensures that the feature heads capture meaningful and prevalent attributes from the data. These components together enable CD to outperform existing methods in scenarios lacking class priors, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
@@
Disentangled representations , Clustering , Contrastive learning , Representation learning , Unsupervised learning
@@ 1 Introduction
The growing availability of large-scale unlabeled datasets has underscored the importance of unsupervised learning in extracting meaningful features without predefined class labels. In computer vision, contrastive learning has become a dominant method for unsupervised representation learning, improving feature extraction by maximizing consistency between augmented views of the same instance while minimizing similarity between distinct instances [1, 2].
Recent contrastive learning methods [3, 4, 5] have incorporated class priors to enhance performance. By using class counts as prior knowledge during network construction, these approaches achieve state-of-the-art results in unsupervised clustering tasks, indicating that class information can lead to more discriminative feature representations. However, this reliance on class priors limits their applicability in real-world scenarios where class information is uncertain.
To address this limitation, we propose Contrastive Disentangling (CD), a novel framework that refines feature extraction and disentanglement in unsupervised learning without relying on class priors. CD integrates multi-level contrastive learning with a disentanglement loss to ensure feature diversity and independence, making it applicable to scenarios lacking class information. Our approach avoids complex additional modules such as pseudo-labeling, focusing instead on the effectiveness of the core framework.
Figure 1: Overview of the Multi-level Contrastive Learning Strategy in CD Framework. The instance-level (top) operates on image instances, while the feature-level (bottom) focuses on features visualized using LIME.
As shown in Figure 1, CD utilizes two levels of contrastive learning: instance-level and feature-level. Instance-level contrastive learning distinguishes image instances in the feature space, while feature-level contrastive learning ensures that distinct prediction heads capture diverse and complementary feature attributes. The feature-level module employs the LIME (Local Interpretable Model-Agnostic Explanations) technique [6] to visualize feature predictions, offering an intuitive interpretation of the model’s behavior. In summary, the main contributions of this paper are:
@@
    Feature Prediction Heads: We introduce feature prediction heads as a novel component in the feature learning process. Through the use of a normalized entropy loss, these heads effectively capture common features prevalent in the dataset, enhancing the representational capacity and generalization of the learned features.
@@
    Feature-Level Contrastive Learning: We propose a feature-level contrastive learning strategy to encourage different feature prediction heads to focus on distinct and complementary features, promoting effective feature disentanglement.
@@
    Simple and Effective Framework: We present a straightforward and effective framework that improves fine-grained feature learning in unsupervised tasks without relying on class priors. Extensive experiments on benchmark datasets demonstrate the superiority of CD, particularly in NMI scores. Ablation studies confirm the effectiveness of each component, providing new insights into unsupervised representation learning.
@@ 2.1 Unsupervised Representation Learning
Unsupervised representation learning seeks to extract meaningful features from unlabeled data. This area is dominated by two main approaches: generative models and contrastive learning. Generative models [7, 8, 9] aim to generate samples that resemble the true data distribution by modeling the underlying latent space. Despite their success in data generation, these models often struggle to learn invariant feature representations—features that remain consistent under various input transformations such as rotation, translation, and scaling [10].
To overcome these limitations, instance-level contrastive learning methods [2, 11, 12, 13] have been developed. These approaches enhance the model’s ability to extract invariant features by maximizing the similarity between different augmented views of the same instance while minimizing the similarity between different instances. As a result, contrastive learning has proven to be highly effective across various downstream tasks, offering robust performance in unsupervised learning.
2.2 Utilization of Class Priors
Several recent approaches [4, 5, 14] have incorporated class prior information to improve model performance by constraining the final outputs based on known class counts. These methods often achieve state-of-the-art results in unsupervised clustering tasks by guiding the model to learn more discriminative feature representations. Additionally, pseudo-labeling techniques are sometimes employed to iteratively refine cluster assignments, further boosting performance.
However, these methods are limited by their reliance on class priors, making them unsuitable for real-world tasks where class labels are often unknown or ambiguous. This reliance highlights the need for methods that can operate effectively without class information.
@@ 2.3 Multi-level Contrastive Learning
Multi-level contrastive learning has emerged as a promising direction within contrastive learning research. The introduction of cluster-level contrastive learning by Li et al. [4] for clustering tasks has opened new avenues for exploring more structured representations. However, methods that partition data based on predefined categories may miss the fine-grained features inherent in more complex datasets, limiting their ability to capture nuanced representations.
Building on this insight, we propose the Contrastive Disentangling (CD) framework, which leverages feature space representations as inputs to multiple feature prediction heads. By combining feature-level and instance-level contrastive learning, CD encourages a decoupling of distinct feature heads, enabling the model to learn more diverse and fine-grained feature representations. This approach advances unsupervised representation learning without relying on class priors, offering a more flexible and robust solution for a variety of tasks.
@@ 3.1 Proposed Contrastive Disentangling Framework
The proposed Contrastive Disentangling (CD) framework leverages a multi-level contrastive learning strategy to achieve fine-grained representation learning without the need for class priors, as illustrated in Figure 2.
Figure 2: Architecture of the proposed Contrastive Disentangling (CD) framework. The upper section illustrates the shared backbone network, instance-level projector, and feature predictor, while the lower section displays the optimization objectives for instance-level and feature-level contrastive learning.
The CD framework comprises the following key components:
@@
    Data Augmentation: Each input batch $\mathbf{x}$ undergoes a series of data augmentation operations, including random cropping, horizontal flipping, color jittering, and Gaussian blurring. These augmentations generate two correlated views, denoted as $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$, of the same sample. This dual-view approach encourages the model to focus on semantically relevant features that are invariant to such transformations.
@@
    Feature Extraction: The augmented views are processed by a shared backbone network $f(\cdot)$, implemented using a ResNet architecture [15]. This backbone extracts feature representations $\mathbf{h}^{(1)}=f(\mathbf{x}^{(1)})$ and $\mathbf{h}^{(2)}=f(\mathbf{x}^{(2)})$, ensuring consistent feature extraction across different views.
@@
    Instance Projector: The extracted features are passed through an instance projector $g(\cdot)$, a multi-layer perceptron (MLP) consisting of two linear layers with batch normalization [16] and ReLU activation functions. This projector maps the features into a latent space, producing latent representations $\mathbf{z}^{(1)}=g(\mathbf{h}^{(1)})$ and $\mathbf{z}^{(2)}=g(\mathbf{h}^{(2)})$. These representations are then used for instance-level contrastive learning.
@@
    Feature Predictor: To capture diverse and disentangled feature attributes, the latent representations are further processed by a feature predictor $h(\cdot)$. This module, structurally similar to the instance projector but with a sigmoid activation in the final layer, outputs predictions $\mathbf{y}^{(1)}=h(\mathbf{z}^{(1)})$ and $\mathbf{y}^{(2)}=h(\mathbf{z}^{(2)})$. Each $\mathbf{y}^{(v)}$ is a matrix of size $N\times K$, where $K$ denotes the number of feature heads. Each feature head is designed to capture a distinct attribute of the data.
@@
    Loss Functions: The total loss function of CD integrates three components: the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, the feature-level contrastive loss $\mathcal{L}_{\text{feat}}$, and the normalized entropy loss $\mathcal{L}_{\text{entropy}}$. These losses work synergistically to promote feature diversity, encourage independence among the learned features, and prevent feature collapse.
@@ 3.2 Instance-level Contrastive Learning
Instance-level contrastive learning aims to learn discriminative features by bringing representations of different augmented views of the same instance closer while pushing apart representations of different instances. This dual objective enhances the model’s ability to distinguish between distinct instances and recognize invariant features across different augmentations.
Given an input batch $\mathbf{x}$ of size $N$, two augmented views $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$ are generated. After passing through the backbone network and instance projector, latent representations $\mathbf{z}^{(1)}$ and $\mathbf{z}^{(2)}$ are obtained. These representations are concatenated to form a combined set of $2N$ latent vectors:
$\mathbf{z}=\begin{bmatrix}\mathbf{z}^{(1)}\\ \mathbf{z}^{(2)}\end{bmatrix}=\begin{bmatrix}\mathbf{z}_{1}\\ \mathbf{z}_{2}\\ \vdots\\ \mathbf{z}_{2N}\end{bmatrix}\in\mathbb{R}^{2N\times d},$ (1)
where $d$ is the dimensionality of the latent space.
The similarity between any two latent representations is measured using cosine similarity:
$\text{sim}_{z}(i,j)=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|\|\mathbf{z}_{j}\|}$ (2)
for $i,j\in\{1,2,\dots,2N\}$. The purpose of using cosine similarity is to measure how closely related two feature vectors are, with a higher value indicating greater similarity. Building on this similarity measure, we define the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, which aims to maximize the similarity between positive pairs (i.e., the other augmented view of the same instance) and minimize the similarity between negative pairs (i.e., views of different instances). For each sample $i$, the instance-level contrastive loss is formulated as:
@@
where $\text{pos}(i)$ denotes the index of the positive sample corresponding to $i$, $\tau_{\text{inst}}$ is a temperature parameter controlling the concentration level of the distribution, and $\mathbbm{1}_{[k\neq i]}$ is an indicator function that equals 1 when $k\neq i$ and 0 otherwise. This formulation follows the NT-Xent loss used in previous work [2, 17, 18, 19]. The total instance-level contrastive loss is computed by averaging over all $2N$ samples:
@@ 3.3 Disentanglement Loss
To ensure that the learned feature representations are both diverse and independent, we introduce two complementary loss functions: the feature-level contrastive loss, which promotes diversity by encouraging each feature head to focus on distinct attributes, and the normalized entropy loss, which ensures that the captured features are meaningful and prevalent across the dataset.
3.3.1 Feature-Level Contrastive Loss
The feature predictor $h(\cdot)$ produces predictions $\mathbf{y}^{(1)}$ and $\mathbf{y}^{(2)}$, each of size $N\times K$, where $K$ is the number of feature heads. To apply feature-level contrastive learning, we concatenate the transposed outputs of the two views:
$\mathbf{y}=\begin{bmatrix}(\mathbf{y}^{(1)})^{\top}\\ (\mathbf{y}^{(2)})^{\top}\end{bmatrix}=\begin{bmatrix}\mathbf{y}_{1}\\ \mathbf{y}_{2}\\ \vdots\\ \mathbf{y}_{2K}\end{bmatrix}\in\mathbb{R}^{2K\times N},$ (5)
where $\mathbf{y}_{i}$ represents the predictions of feature head $i$ across the batch. The cosine similarity between any two feature prediction vectors is defined as:
$\text{sim}_{y}(i,j)=\frac{\mathbf{y}_{i}\cdot\mathbf{y}_{j}}{\|\mathbf{y}_{i}\|\|\mathbf{y}_{j}\|}$ (6)
The feature-level contrastive loss encourages each feature head to capture distinct attributes by maximizing the similarity between corresponding heads across different views while minimizing the similarity between different heads. For feature head $i$, the loss is defined as:
@@ 3.3.1 Feature-Level Contrastive Loss
where $\text{pos}(i)$ denotes the index of the positive sample for feature head $i$, and $\tau_{\text{feat}}$ is the temperature parameter for the feature-level contrastive loss. The total feature-level contrastive loss is computed by averaging over all $2K$ feature heads:
@@
3.3.2 Normalized Entropy Loss
To prevent feature collapse and ensure that the learned features are meaningful and widespread across the dataset, we introduce a normalized entropy loss $\mathcal{L}_{\text{entropy}}$. This loss encourages each feature head to distribute its predictions evenly, capturing non-trivial and prevalent patterns in the data. The normalized binary entropy for each feature prediction vector $\mathbf{y}_{i}$ is defined as:
$H(\mathbf{y}_{i})=-\frac{1}{N\log(2)}\sum_{j=1}^{N}\left(\mathbf{y}_{ij}\log(\mathbf{y}_{ij})+(1-\mathbf{y}_{ij})\log(1-\mathbf{y}_{ij})\right)$ (9)
The overall normalized entropy loss is the average entropy across all feature heads:
$\mathcal{L}_{\text{entropy}}=\frac{1}{2K}\sum_{i=1}^{2K}H(\mathbf{y}_{i})$ (10)
3.3.3 Combined Disentanglement Loss
The total disentanglement loss is the combination of the feature-level contrastive loss and the normalized entropy loss:
$\mathcal{L}_{\text{disentanglement}}=\mathcal{L}_{\text{feat}}-\alpha\mathcal{L}_{\text{entropy}}$ (11)
where $\alpha$ is a hyperparameter that controls the balance between encouraging diversity and ensuring meaningful feature representations. By jointly optimizing these two objectives, the feature heads are encouraged to capture both diverse and informative patterns, leading to higher-quality feature disentanglement.
3.4 Total Loss
The total loss function for training the CD framework is the sum of the instance-level contrastive loss and the disentanglement loss. This ensures that both the instance-level discrimination and feature-level disentanglement are optimized simultaneously:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{inst}}+\mathcal{L}_{\text{disentanglement}}.$ (12)
By optimizing both the instance and feature-level objectives, our framework achieves a robust balance between learning discriminative instance-level representations and promoting diverse, meaningful feature representations without relying on class priors. Algorithm 1 summarizes the proposed method.